OpenStudy (anonymous):

A circle has a diameter of 12 inches find the area of the circle use 3.14

OpenStudy (anonymous):

diameter is twice the radius and the area of a circle is pi times radius squared so 3.14 x 6^2 sq inches

OpenStudy (anonymous):

Radius=12/2=6 \[Area=\Pi*r ^{2}\] =3.14*6^2=113.04

OpenStudy (anonymous):

11304?

OpenStudy (anonymous):

No. See the decimal point?

OpenStudy (anonymous):

Its 113.04

OpenStudy (anonymous):

113.04

OpenStudy (anonymous):

Right

OpenStudy (anonymous):

sq.inches. it is important to note that the answer is in square inches, otherwise it wouldn't make any sense.